OpenStudy (anonymous):

Pat owns a factory that manufactures key chains. His weekly profit (in hundreds of dollars) is given by P(x)= -2x^2+60x-120, where x is the number of key chains sold. a) what is the largest number of cases he can sell and still make a profit? b) how many cases should he make in sell in order to maximize his profits?

OpenStudy (danjs):

how many in a case of em

OpenStudy (danjs):

question / answer mismatch thing

hartnn (hartnn):

for a) we need to have a positive profit, so can you solve \( -2x^2+60x-120 > 0 \) that will give the largest number of "keys". Once we know how many keys are there in a case, we can find the largest number of cases. Please provide all the required information :)

OpenStudy (anonymous):

This is why I am confused on this problem..that is the whole problem word for word...The answer for A is 27 and B is 15 I just don't understand how they get B.

hartnn (hartnn):

ok, that means we assume there is one key in one case!

hartnn (hartnn):

start by solving \(-2x^2+60x-120 > 0\) need a helping hand with that?

OpenStudy (anonymous):

\[-60+/-\sqrt{(60^2-4(-2)(-120)} \over 2(-2)\]

hartnn (hartnn):

ohh, just saw you need help with only b) ? to maximize the profit, we need to find the extrema of P(x). that we do by finding the derivative and setting it to 0. P'(x) = 0 \(\dfrac{d}{dx}[−2x^2+60x−120] = 0 \)

hartnn (hartnn):

the easier way to approach \(−2x^2+60x−120 > 0\) is to factor out -2 first! \(-2 (x^2 - 30 x + 60) > 0\) now using the quadratic formula for \(x^2 -30x +60\) would be easier, right? :)

OpenStudy (anonymous):

yep, my bad!

hartnn (hartnn):

go ahead with the solution, let me know if you get stuck...have you already got 27 for a) ?

OpenStudy (anonymous):

yes I did, just not sure what the derivative is.

OpenStudy (anonymous):

or is 27 the derivative

hartnn (hartnn):

ohh... you've not been taught derivatives yet? have you been taught how to find the vertex of \(ax^2+bx+c\) ??

OpenStudy (anonymous):

yes -b/2*a

hartnn (hartnn):

excellent! so the x co-ordinate of vertex of x^2-30x+60 will give you the max profit!

OpenStudy (anonymous):

thank you so much!

hartnn (hartnn):

welcome ^_^

OpenStudy (astrophysics):

First thing that I thought of was the vertex haha